# Data-Juicer Sandbox: A Feedback-Driven Suite for Multimodal Data-Model Co-development

### v2

@@ 4.2 Ranking Single-Operator Data Pools
Observation 1 (Data vs. Model) Multimodal models’ efficacy is closely tied to the fidelity of their output modalities, which can be explicitly reflected in filtering of input training data.
For text-to-video task, all top-3 OPs are video-only OPs. For image-to-text and image-text pre-training tasks, two of their top-3 OPs are text-only, and image-text OPs respectively. This trend of influential OPs holds beyond top-3 ranks (detailed in Appendix E.2), suggesting that more attention and resources should be allocated to data processing related to the output modalities of the studied models.
@@
Observation 2 (Diversity vs. Quality) In contrast, data diversity is more crucial for image-to-text model, while data quality is key for text-to-video and image-text pre-training models.
@@ 4.3 Shaping Data Recipes of Top OP Combinations
Although the Top-3 OP recipes exhibit suboptimal performance, we observe positive gains when combining some pairs of OPs in them can outperform both single top-1 and top-3 combinations, such as in the case of image-to-text task when combining TextActionFilter and LanguageIDScoreFilter.
4.4 Scaling Data Samples and Computes
@@
Observation 7 (Effect of Compute Scaling) Scaling up compute resources with high-quality data benefits both the image-to-text, text-to-video, and image-text pre-training tasks. Among these, the image-text pre-training task demonstrates clear scaling law behaviors in exponential form.
@@
Moreover, for the image-text pre-training task, Fig. LABEL:fig:scaling-lines illustrates improvements as both model size and compute scale increase. Notably, the linear growth in relative improvement with the exponential increase in computation aligns with known scaling laws (Cherti et al., 2023), suggesting that expanding model size and compute resources, alongside processing data via our high-quality recipes, yields consistent scaling of performance gains.
4.5 Transferring Recipes to Larger-Scale Scenarios
The previous image-text pre-training experiments on different CLIP models and compute resources (Fig. LABEL:fig:scaling-lines) have shown good potential to scale up the benefits from data processed by the top recipes. In this section, we further apply the gained insights and top data recipes into larger-scale models and datasets for the other two image-to-text and text-to-video tasks.
@@ 4.5 Transferring Recipes to Larger-Scale Scenarios
Table 2 summarizes the results. Since the official MGM-2B model has not been evaluated on MMBench-CN, we trained a reproduced version using the official training scripts, and its performance is closely aligned with the benchmarks reported for the official model. We can see that compared to the baseline trained on the full dataset, our model achieves superior performance while trained on only 1/10 of the distinct instances and 1/2 of its original total instances (i.e., using half of baseline’s compute FLOPs). Notably, they differ only in the processing of the pre-training data, with the model and training configurations being consistent, underscoring the effectiveness of our insights.
@@
Table 3 showcases our notable performance on VBench leaderboard, reporting average scores of 16 metrics from both quality and semantic dimensions. We first enhance T2V-Turbo (the last row) with data-enhanced distillation training on 147k instances (the second row), and then self-distill it with other 228k instances (the first row). Compared to the base model in last row, our method yields notable uplifts of 1.53% and 2.59% on the Board and Uniform Average score respectively. This verifies the effectiveness of our sandbox and insights again, which leverages VideoFramesTextSimilarityFilter OP to enhance the video-prompt alignment, and the VideoNSFWFilter OP to ensure high quality of the generated video.
Also note that these results highlight the data-efficiency of our method to reduce overall cost with dedicated data-model co-development. Taking the third-place model, VEhancer, as an example, it mainly focuses on architectural refinements building upon the foundation of VC2. While our model also originated from VC2, it gains a higher performance boost but with at least 22x less compute costs than VEhancer. More detailed calculation of FLOPs is presented in Appendix  E.9. Besides, in Appendix E.7, we further conduct ablation study to verify the effectiveness of our data-model co-improvement over the base model T2V-Turbo, including aspects from data (high-quality dataset with our derived recipes), model (LoRA initialization and self-distillation), and data-model co-design (real-data loss and self-evolution). These results demonstrate the potential cost-effectiveness of the proposed sandbox when utilizing fine-grained feedback from both the data and the model simultaneously.
@@ 5 Conclusion and Future Works
Looking ahead, we will continuously extend the sandbox’s compatibility to encompass more model-centric infrastructures and develop more off-the-shelf and effect-proven workflows, to facilitate reuse and expedite innovation. Besides, due to the resource limitation, we will explore transferring the insights in larger-scale CLIP models beyond our experiments, and theoretically investigating the trade-off between sandbox cost and feedback transferability, such as taking the sampling ratio of data pools as a tunable knob.
@@
Our appendix is organized as follows:
In Sec. A, we present a more detailed introduction to related works.
In Sec. B, we present more details on the proposed sandbox suite from an infrastructure perspective, including the overview architecture (Sec.B.1), the capability factory and behavior hooks (Sec. B.2) and its extensibility (Sec. B.3).
In Sec. C, we present detailed analysis and discussion on the cost of the proposed sandbox workflow.
In Sec. D, we present implementation details about the sandbox experiments, including
  • Sec. D.1: the two strategies used to combine multi-OPs into recipes based on their correlations.
  • Sec. D.2: the adopted model evaluation metrics.
  • Sec. D.3: the functionalities and corresponding statistics of the studied Data-Juicer OPs.
  • Sec. D.4: the detailed settings about the image-to-text use case.
  • Sec. D.5: the detailed settings about the text-to-video use case.
  • Sec. D.6: the detailed settings about the image-text pre-training use case.
In Sec. E, we present more experimental results, including
  • Sec. E.1: the summarization of the studied scope to vary in the sandbox experiments, in terms of main effectiveness evidence, model types, dataset size, and compute scale.
  • Sec. E.2: the complete operator ranking in single-OP experiments.
  • Sec. E.3: the data diversity analysis in multiple-OP experiment.
  • Sec. E.4: the correlation analysis in multiple-OP experiment.
  • Sec. E.5: performance of models trained with recipes derived from the correlation analysis.
  • Sec. E.6: the detailed experimental results for Image-to-Text task.
  • Sec. E.7: detailed ablation study of the Data-Juicer T2V models, from the view of data-model co-improvements over its base teacher T2V model, including the view of data (high-quality dataset with our derived recipes), model (LoRA initialization and self-distillation) and data-model co-design (real-data loss and self-evolution).
  • Sec. E.8: the full V-Bench leaderboard results.
  • Sec. E.9: the detailed calculation of the reported FLOPs.
@@ Appendix A Discussion on More Related Works
Multimodal large models have captivated researchers with their formidable capabilities (OpenAI, 2024a, b), leading to a surge in model-centric development efforts. These focuses mainly lie in refining training algorithms (Caffagni et al., 2024; Li et al., 2024a; Zhang et al., 2024), advancing model architectures and components (He et al., 2024a; Yin et al., 2024; Jiao et al., 2024b), and harnessing the models’ potential for various applications (Wang et al., 2024a; Liu et al., 2024; Zhou et al., 2024a). There is a growing consensus that transformer-based scaling is predominant (Xu et al., 2023). However, the high computational requirements imposed by scaling laws (Xu et al., 2024c) and the optimization challenges inherent to large models (Manduchi et al., 2024) often confine insights to specific datasets or vague data characteristics. This situation leaves a significant gap in comprehending the extent to which models’ performance and behavior hinge upon implicit assumptions and inductive biases embedded within the underlying data distributions. In contrast, our work demonstrates a feasible and promising path to fill in this gap by explicitly linking data processing effects with the downstream performance of trained models through numerous contrastive sandbox experiments.
@@
An emerging trend shifts the focal point from models to data (Jakubik et al., 2024; Bai et al., 2024; Chen et al., 2024c; Jiao et al., 2024a), underscored by the notion that large models function akin to data compressors (Delétang et al., 2024). Echoing the principle of “garbage in, garbage out”, meticulous data processing is recognized as pivotal. Efforts now isolate data manipulation as a primary experimental variable in multimodal generative modeling (He et al., 2024b). Nonetheless, multimodal data processing involves highly heterogeneous processing workflows, vast quantities, diverse types, and the high cost of training downstream models. This complexity results in predominantly heuristic approaches, such as data filtering and synthesis guided by human intuition (Long et al., 2024; Zhou et al., 2024b).
@@
The landscape for multimodal model development has advanced significantly, boasting a variety of strong infrastructures and frameworks for training and evaluation. Prominent examples include Transformers (Wolf et al., 2020), Diffusers (von Platen et al., 2022), NeMo (Harper et al., ), MMagic (MMagic Contributors, 2023), ESPNet (Peng et al., 2023), and MMBench (Liu et al., 2023).
However, when it comes to multimodal data infrastructure, the primary contributions have been datasets and dataset-specific preprocessing tools, such as DatasetsHub from HuggingFace (Lhoest et al., 2021). The standardization and efficient utilization of practical expertise and foundational data processing capabilities remain unaddressed. Existing frameworks for data processing predominantly focus on single-modal data (TogetherAI, 2023; Bradski, 2000; Hwang et al., 2023), underlining the early stages of development in systematic platforms for multimodal data (Chen et al., 2024a; Du et al., 2023; Chen et al., 2024b).
@@
Appendix B Details to the Sandbox Suite from Infrastructure Perspective
@@ B.1 Overview Architecture
Figure 7: Overview of the Data-Juicer Sandbox Laboratory. The workflow involves four stages, each allowing flexible customization at different levels within the data-model co-development lifecycle. More details can be found in Section B
To support co-development based on feedback from data and model, we design the layered sandbox laboratory as illustrated in Figure 7. The laboratory incorporates a spectrum of components for activities such as data analysis, filtering, recipe optimization, model training, inference, and evaluation, all orchestratable via configuration files. The architecture is stratified into three tiers: bespoke end-to-end workflows, generic development behaviors, and foundational data-model development capabilities.
@@
This design allows for streamlined configuration and reuse of established infrastructure. Importantly, it expedites the prototyping of data and model development solutions, integrating actionable and measurable capabilities for swift feedback derivation and informed decision-making. Concrete explorations of the lower tiers are provided in Section B.2, while Section 3.2 showcases the sandbox’s utility in establishing a probe-analyze-refine workflow for data-model co-development.
@@ B.2 Flexible Capability Factory and Behavior Hooks
The sandbox also provides many classes to encapsulate observational capabilities, enabling subsequent optimization. For example, Data-Juicer-based classes can efficiently compute metrics such as text perplexity, video aesthetic value, and image quality scores with GPT API calls, along with statistical values like mean, variance, and percentiles. For models, various evaluation benchmarks are supported, like VBench and FVD (Unterthiner et al., 2019) for synthesized video assessment, and TextVQA (Singh et al., 2019), MMBench (Liu et al., 2023), and MME (Fu et al., 2023) for image-to-text evaluation.
@@
python tools/sandbox_starter.py --config configs/demo/sandbox.yaml
@@ B.3 Extensibility of the Sandbox Suite
Appendix C Analysis on the Costs of the Proposed Sandbox Workflow
@@ D.1 Strategies to Combine OPs According to Correlations
In addition to assembling the OPs with the overall best performance, we also incorporate an analysis of inter-OP relationships into our recipe formulation process. Our workflow accommodates two strategies, with specific applications detailed in Section E.4:
@@ D.2 Evaluation Metrics
TextVQA, MMBench, MME. These benchmarks serve as critical evaluators of MLLM’s proficiency in understanding images. TextVQA (Singh et al., 2019) specifically targets the assessment of MLLMs’ abilities to read and reason about textual content embedded within images. MMBench (Liu et al., 2023), a vast multimodal benchmark, encompasses perception and reasoning skills through a plethora of multi-choice questions, numbering in the thousands. Additionally, a Chinese translation, MMBench-CN, is integrated for broader accessibility. MME (Fu et al., 2023) focuses on the perceptual and cognitive competencies of MLLMs, incorporating 14 finely categorized subtasks, each addressing Yes/No inquiries underpinned by meticulously crafted guidelines.
@@ D.4 Image-to-Text Generation
We first conduct single-OP experiments (Section 4.2) that encompasses 22 text-image relevant OPs from Data-Juicer, split evenly between text-only and image-related multimodal OPs. After the two-stage training, model evaluation is conducted on established benchmarks including TextVQA (Singh et al., 2019), MMBench (Liu et al., 2023), and MME (Fu et al., 2023).
For multi-OP data pools (Section 4.3), we identify the top-3 highest-performing OPs from single-OP experiments and study their possible combinations. Additionally, we analyze the correlations among the 23 data statistics produced by 22 OPs capable of generating instance-level stats  [^2] The image height and width are produced by one OP.. Employing a hierarchical clustering algorithm  (Ward Jr, 1963), these OPs are grouped into three clusters based on correlation coefficients, with the highest-performing OP from each cluster selected for combination testing. To ensure a robust and fair comparison, we must acknowledge the constraints imposed by the limited data volume within the highest-tier data pool. As the number of combinations increases, the available dataset size diminishes. In particular, the size of $\mathcal{P}_{\mathcal{S}}$ was reduced from 200k samples to 159k samples during the Top-3 combination experiments. Similarly, in the cluster-wise combination experiments, the dataset size decreased from 200k samples to 126k samples.
Next, in Section 4.4, we explore the optimal OP combination based on previous experiments and adopt the methodology from Section 3.2.3 for comparative experiments on training with repeated data versus non-repeated data. Note that due to filtering, the final instance count decreases from 200k to approximately 159k after the OP combination. These data are then repeated in increments from double to eightfold, mirroring the size of the original pretraining set.
Collectively, all these experiments yield profound insights into image-to-text model training, data processing, and iteration strategies from a data-model co-development perspective, further verified in the larger-scale scenario in Section 4.5.

### v3

@@ 4.2 Ranking Single-Operator Data Pools
Observation 1 (Data vs. Model) Multimodal models’ efficacy is closely tied to the fidelity of their output modalities, which can be explicitly reflected in the filtering of input training data.
For the text-to-video task, all top-3 OPs are video-only OPs. For image-to-text and image-text pre-training tasks, two of their top-3 OPs are text-only, and image-text OPs respectively. This trend of influential OPs holds beyond top-3 ranks (detailed in Appendix E.2), suggesting that more attention and resources should be allocated to data processing related to the output modalities of the studied models.
@@
Observation 2 (Diversity vs. Quality) In contrast, data diversity is more crucial for image-to-text models, while data quality is key for text-to-video and image-text pre-training models.
@@ 4.3 Shaping Data Recipes of Top OP Combinations
Although the Top-3 OP recipes exhibit suboptimal performance, we observe positive gains when combining some pairs of OPs in them can outperform both single top-1 and top-3 combinations, such as in the case of the image-to-text task when combining TextActionFilter and LanguageIDScoreFilter.
4.4 Scaling Data Samples, Computes & Model Size
@@
Observation 7 (Effect of Compute Scaling) Scaling up compute resources with high-quality data benefits both the image-to-text, text-to-video, image-text pre-training, and image-captioning tasks. Among these, the image-text pre-training and image-captioning tasks demonstrate clear scaling law behaviors in exponential form across different dataset and model sizes.
@@
For the image-text pre-training task, Fig. LABEL:fig:scaling-lines illustrates improvements as both model size and compute scale increase. Notably, the linear growth in relative improvement with the exponential increase in computation aligns with known scaling laws (Cherti et al., 2023), suggesting that expanding model size and compute resources, alongside processing data via our high-quality recipes, yields consistent scaling of performance gains.
Moreover, we conduct experiments using InternVL2.0 across various scales (1B, 2B, 4B, 26B parameters), with the results summarized in Fig. LABEL:fig:scaling-internvl (x-axis displayed on a log scale). We initially experimented with the smallest model (1B parameters) on 23 single-OP and 6 multi-OP data pools to identify the optimal data recipe. The top-3 combination recipes derived from this initial phase were subsequently applied to all selected model scales. Our findings indicate that all three recipes consistently maintain significant performance advantages as the model scale increases from 1B to 26B parameters, demonstrating clear scaling law behaviors.
@@
4.5 Transferring Recipes to Larger-Scale Heterogeneous Scenarios
The previous image-text pre-training and image-caption fine-training experiments, conducted on CLIP and InternVL models of varying sizes (Fig. LABEL:fig:scaling-lines and Fig. LABEL:fig:scaling-internvl), already shown that top-performing recipes identified at small scales can be effectively utilized when compute resources are scaled up. In this section, we further apply the gained insights and top data recipes to larger-scale heterogeneous models (i.e., across different model architectures) and heterogeneous datasets (i.e., from diverse data sources) for the other two tasks (image-to-text and text-to-video).
@@ 4.5 Transferring Recipes to Larger-Scale Heterogeneous Scenarios
Table 2 summarizes the results. Since the official MGM-2B model has not been evaluated on MMBench-CN, we trained a reproduced version using the official training scripts, and its performance is closely aligned with the benchmarks reported for the official model. We can see that compared to the baseline trained on the full dataset, our model achieves superior performance while trained on only 1/10 of the distinct instances and 1/2 of its original total instances (i.e., using half of the baseline’s compute FLOPs). Notably, they differ only in the processing of the pre-training data, with the model and training configurations being consistent, underscoring the effectiveness of our insights.
@@
Table 3 showcases our notable performance on the VBench leaderboard, reporting average scores of 16 metrics from both quality and semantic dimensions. We first enhance T2V-Turbo (the last row) with data-enhanced distillation training on 147k instances (the second row), and then self-distill it with other 228k instances (the first row). Compared to the base model in the last row, our method yields notable uplifts of 1.53% and 2.59% on the Board and Uniform Average scores respectively. This verifies the effectiveness of our sandbox and insights again, which leverages VideoFramesTextSimilarityFilter OP to enhance the video-prompt alignment, and the VideoNSFWFilter OP to ensure the high quality of the generated video.
Note that these results highlight the data-efficiency of our method to reduce overall cost with dedicated data-model co-development. Taking the third-place model, VEhancer, as an example, it mainly focuses on architectural refinements building upon the foundation of VC2. While our model also originated from VC2, it gains a higher performance boost but with at least 22x less compute costs than VEhancer. Detailed calculation of FLOPs is presented in Appendix  E.9. Besides, in Appendix E.7, we further conduct an ablation study to verify the effectiveness of our data-model co-improvement over the base model T2V-Turbo, including aspects from data (high-quality dataset with our derived recipes), model (LoRA initialization and self-distillation), and data-model co-design (real-data loss and self-evolution). These results demonstrate the potential cost-effectiveness of the proposed sandbox when utilizing fine-grained feedback from both the data and the model simultaneously.
4.6 Additional Experiment Results and Details
Further Analysis and Ablation Studies. Appendix E.3 and E.4 detail the data diversity and correlation analyses on our multiple-OP experiments, respectively. They showcase the analysis ability provided by the suite with word entropy, word cloud visualization, and visualizations of Pearson correlation coefficients among OP statistics and benchmark performance. Appendix E.5 evaluates the performance of models trained using recipes derived from the correlation analysis. Appendix E.7 provides a detailed ablation study of the Data-Juicer T2V models, focusing on data-model co-improvements beyond the base teacher T2V model. This includes analyses from the perspective of data (high-quality datasets with derived recipes), model (LoRA initialization and self-distillation), and data-model co-design (real-data loss and self-evolution).
[…]
Complete Results. As for the complete results of other previously mentioned experiments. Appendix E.2 presents the complete operator ranking from single-operator experiments. Appendix E.6 provides comprehensive results for the Image-to-Text task. The full V-Bench leaderboard results are listed in Appendix E.8. Finally, Appendix E.9 details the calculation of reported FLOPs.
Implementation Details. We provide a summary of the varying scope of our sandbox experiments in Appendix E.1, detailing aspects such as main effectiveness evidence, model types, dataset size, and compute scale. Besides, we present a discussion in Appendix D.1 on the two strategies employed to combine multiple operators into recipes based on their correlations. Appendix D.2 outlines the model evaluation metrics adopted. The functionalities and corresponding statistics of the studied Data-Juicer operators are detailed in Appendix D.3. Furthermore, Appendix D.4 elaborates on the detailed settings for the image-to-text use case, Appendix D.5 for the text-to-video use case, and Appendix D.6 for the image-text pre-training use case.
@@ 5 Conclusion and Future Works
To facilitate reuse and expedite innovation, we will continuously extend the sandbox’s compatibility to encompass more model-centric infrastructures and develop more off-the-shelf and effect-proven workflows, such as for reinforced fine-tuning (Pan et al., 2025). Besides, we will theoretically investigating the trade-off between sandbox cost and feedback transferability, such as taking the sampling ratio of data pools as a tunable knob.
@@
[…]
1Table of Contents
@@ Appendix A Discussion on More Related Works
Multimodal large models have captivated researchers with their formidable capabilities (OpenAI, 2024a, b), leading to a surge in model-centric development efforts. These focuses mainly lie in refining training algorithms (Caffagni et al., 2024; Li et al., 2024a; Zhang et al., 2024), advancing model architectures and components (He et al., 2024a; Yin et al., 2024; Jiao et al., 2024), and harnessing the models’ potential for various applications (Wang et al., 2024a; Liu et al., 2024a; Zhou et al., 2024a). There is a growing consensus that transformer-based scaling is predominant (Xu et al., 2023). However, the high computational requirements imposed by scaling laws (Xu et al., 2024c) and the optimization challenges inherent to large models (Manduchi et al., 2024) often confine insights to specific datasets or vague data characteristics. This situation leaves a significant gap in comprehending the extent to which models’ performance and behavior hinge upon implicit assumptions and inductive biases embedded within the underlying data distributions. In contrast, our work demonstrates a feasible and promising path to fill in this gap by explicitly linking data processing effects with the downstream performance of trained models through numerous contrastive sandbox experiments.
@@
An emerging trend shifts the focal point from models to data (Jakubik et al., 2024; Bai et al., 2024; Chen et al., 2024c; Jiao et al., 2025), underscored by the notion that large models function akin to data compressors (Delétang et al., 2024). Echoing the principle of “garbage in, garbage out”, meticulous data processing is recognized as pivotal. Efforts now isolate data manipulation as a primary experimental variable in multimodal generative modeling (He et al., 2024b). Nonetheless, multimodal data processing involves highly heterogeneous processing workflows, vast quantities, diverse types, and the high cost of training downstream models. This complexity results in predominantly heuristic approaches, such as data filtering and synthesis guided by human intuition (Long et al., 2024; Zhou et al., 2024b).
Figure 9: Overview of the Data-Juicer Sandbox Laboratory. The workflow involves four stages, each allowing flexible customization at different levels within the data-model co-development lifecycle. More details can be found in Sec. B
@@
The landscape for multimodal model development has advanced significantly, boasting a variety of strong infrastructures and frameworks for training and evaluation. Prominent examples include Transformers (Wolf et al., 2020), Diffusers (von Platen et al., 2022), NeMo (Harper et al., ), MMagic (MMagic Contributors, 2023), ESPNet (Peng et al., 2023), and MMBench (Liu et al., 2024b).
However, when it comes to multimodal data infrastructure, the primary contributions have been datasets and dataset-specific preprocessing tools, such as DatasetsHub from HuggingFace (Lhoest et al., 2021). The standardization and efficient utilization of practical expertise and foundational data processing capabilities remain unaddressed. Existing frameworks for data processing predominantly focus on single-modal data (Weber et al., 2024; Bradski, 2000; Hwang et al., 2023), underlining the early stages of development in systematic platforms for multimodal data (Chen et al., 2024a; Du et al., 2023; Chen et al., 2024b).
@@
Appendix B Sandbox Suite from Infrastructure Perspective
@@ B.1 Overview Architecture
To support co-development based on feedback from data and model, we design the layered sandbox laboratory as illustrated in Fig. 9. The laboratory incorporates a spectrum of components for activities such as data analysis, filtering, recipe optimization, model training, inference, and evaluation, all orchestratable via configuration files. The architecture is stratified into three tiers: bespoke end-to-end workflows, generic development behaviors, and foundational data-model development capabilities.
@@
This design allows for streamlined configuration and reuse of established infrastructure. Importantly, it expedites the prototyping of data and model development solutions, integrating actionable and measurable capabilities for swift feedback derivation and informed decision-making. Concrete explorations of the lower tiers are provided in Sec. B.2, while Sec. 3.2 showcases the sandbox’s utility in establishing a probe-analyze-refine workflow for data-model co-development.
@@ B.2 Flexible Capability Factory and Behavior Hooks
The sandbox also provides many classes to encapsulate observational capabilities, enabling subsequent optimization. For example, Data-Juicer-based classes can efficiently compute metrics such as text perplexity, video aesthetic value, and image quality scores with GPT API calls, along with statistical values like mean, variance, and percentiles. For models, various evaluation benchmarks are supported, like VBench and FVD (Unterthiner et al., 2019) for synthesized video assessment, and TextVQA (Singh et al., 2019), MMBench (Liu et al., 2024b), and MME (Fu et al., 2023) for image-to-text evaluation.
@@
{minted}
[fontsize=, bgcolor=lightblue]bash python tools/sandbox_starter.py  –config configs/demo/sandbox.yaml
@@ B.3 Extensibility of the Sandbox Suite
Appendix C Analysis on the Costs of the Proposed Workflow
@@ D.1 Strategies to Combine OPs According to Correlations
In addition to assembling the OPs with the overall best performance, we also incorporate an analysis of inter-OP relationships into our recipe formulation process. Our workflow accommodates two strategies, with specific applications detailed in Sec. E.4:
@@ D.2 Evaluation Metrics
TextVQA, MMBench, MME. These benchmarks serve as critical evaluators of MLLM’s proficiency in understanding images. TextVQA (Singh et al., 2019) specifically targets the assessment of MLLMs’ abilities to read and reason about textual content embedded within images. MMBench (Liu et al., 2024b), a vast multimodal benchmark, encompasses perception and reasoning skills through a plethora of multi-choice questions, numbering in the thousands. Additionally, a Chinese translation, MMBench-CN, is integrated for broader accessibility. MME (Fu et al., 2023) focuses on the perceptual and cognitive competencies of MLLMs, incorporating 14 finely categorized subtasks, each addressing Yes/No inquiries underpinned by meticulously crafted guidelines.
@@ D.4 Image-to-Text Generation
We first conduct single-OP experiments (Sec. 4.2) that encompasses 22 text-image relevant OPs from Data-Juicer, split evenly between text-only and image-related multimodal OPs. After the two-stage training, model evaluation is conducted on established benchmarks including TextVQA (Singh et al., 2019), MMBench (Liu et al., 2024b), and MME (Fu et al., 2023).
For multi-OP data pools (Sec. 4.3), we identify the top-3 highest-performing OPs from single-OP experiments and study their possible combinations. Additionally, we analyze the correlations among the 23 data statistics produced by 22 OPs capable of generating instance-level stats  [^2] The image height and width are produced by one OP.. Employing a hierarchical clustering algorithm  (Ward Jr, 1963), these OPs are grouped into three clusters based on correlation coefficients, with the highest-performing OP from each cluster selected for combination testing. To ensure a robust and fair comparison, we must acknowledge the constraints imposed by the limited data volume within the highest-tier data pool. As the number of combinations increases, the available dataset size diminishes. In particular, the size of $\mathcal{P}_{\mathcal{S}}$ was reduced from 200k samples to 159k samples during the Top-3 combination experiments. Similarly, in the cluster-wise combination experiments, the dataset size decreased from 200k samples to 126k samples.
Next, in Sec. 4.4, we explore the optimal OP combination based on previous experiments and adopt the methodology from Sec. 3.2.3 for comparative experiments on training with repeated data versus non-repeated data. Note that due to filtering, the final instance count decreases from 200k to approximately 159k after the OP combination. These data are then repeated in increments from double to eightfold, mirroring the size of the original pretraining set.
Collectively, all these experiments yield profound insights into image-to-text model training, data processing, and iteration strategies from a data-model co-development perspective, further verified in the larger-scale scenario in Sec. 4.5.